OpenStudy (anonymous):

Can anyone solve this problem? P=2L +2w Solve for L

OpenStudy (anonymous):

L=\[(P-2w)/2\]

OpenStudy (anonymous):

ok so p-2w=2l .5p-w=l

OpenStudy (anonymous):

P-2W=2L L=(P-2W)/2

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

L=(P-2W)\2

OpenStudy (anonymous):

durga.kulla, why is it 5p?